a cable car of a mass 12 kg. is parked on a slope of 20° compressing a giant retaining spring to a length of 0.75m. if the spring constant is 2.0×10⁴n/m. what is the length of the spring when relaxed? neglect friction
1
Expert's answer
2019-03-25T09:09:12-0400

For the equilibrium:


"mg \sin{20°} =k(L-l)"

"12(9.8) \sin{20°} =20000(L-0.75)"

"L= 0.7520 m."